OpenStudy (anonymous):

(1-x^2) dy/dx - xy = 1

OpenStudy (anonymous):

i bet this question should be answered by using linear, somehow, I get stuck on halfway.

OpenStudy (anonymous):

yup it is linear differential eqn did u find integrating factor? If yes then what did u get?

OpenStudy (anonymous):

i did it, somehow, this is where i stuck, wait, need to draw it first @.@

OpenStudy (anonymous):

i guess write it would be better so, 1stly, i divide whole eqn with (1-x^2)

OpenStudy (anonymous):

I.F. = exp{∫ -x/(1-x^2) dx } put 1-x^2 = t -2xdx= dt after plugging the values i m getting sqrt(t) or sqrt (1-x^2)

OpenStudy (anonymous):

turn out like this dy/dx - (xy/1-x^2) = 1/(1-x^2) then, my p(x)=(-x/1-x^2) q(x)=1/(1-x^2) then, I integrated p(x) and ended up /p(x)= ...

OpenStudy (anonymous):

oh, i need to subtitute it ? D:

OpenStudy (anonymous):

yeah, that's what i got before, then, that's the part where i stuck

OpenStudy (anonymous):

i multiply the whole eqn with surd(1-x^2)

OpenStudy (anonymous):

but im not sure if im in the right path or not, so, it's right isn't ?

OpenStudy (anonymous):

if that's correct, i shall proceed with next step :D

OpenStudy (anonymous):

yah u r going right u need to multiply the I.F on both sides of the equation

OpenStudy (anonymous):

and this is what i got

OpenStudy (anonymous):

\[y \sqrt{1-x^{2}}=\int\limits_{?}^{?}(1/(1-x^{2}))\]

OpenStudy (anonymous):

so, am i in the right track ?

OpenStudy (anonymous):

LHS u r right but in RHS u will get ∫ 1/ sqrt (1-x^2) dx

OpenStudy (anonymous):

ehh, wait, i forgot to multiply it with sqrt (1-x^2) but as far as i know, my RHS suppose to be q(x) times I.F right ? so, isn't it will turn out like this :- q(x) = 1/(1-x^2) I.F = sqrt (1-x^2) so, RHS = sqrt (1-x^2)/(1-x^2)

OpenStudy (anonymous):

yup u r proceeding in the right direction

OpenStudy (anonymous):

oh, oke, i'll proceed :D

OpenStudy (anonymous):

but ... lol, im kinda stuck over here lets say that i subtitute t=1-x^2 so i have my RHS = (t^1/2)/t RHS = t^(1/2 - 1) RHS = t^(-1/2) am I right ? :D

OpenStudy (anonymous):

so, i've to replace it back and turn to RHS = 1/squareroot(1-x^2)

OpenStudy (anonymous):

y r u doing substitution it's a direct formula ∫ 1/ sqrt(1-x^2) dx = sin(x)

OpenStudy (anonymous):

O.o i forgot bout its existence

OpenStudy (anonymous):

soo, this is where I wrong, haha, finally got the solution, thanks a lot ! :D

OpenStudy (anonymous):

cool :-D

OpenStudy (anonymous):

btw, how did u type "∫" ? *sry, im pretty new here*